# Graphical lasso for extremes

### v2

@@ 1 Introduction
In this paper, we propose a direct estimate of $\Theta$ with a built-in option for sparse estimation via $L_{1}$-regularization. We term it the extreme graphical lasso. The core idea is as follows. We show that by adding a positive constant $c$ to each entry of $\Theta$, the matrix
@@
where $\gamma>0$ is the regularization parameter. The extreme graphical lasso requires solving only one optimization problem and therefore is efficient in handling large dimensions. In addition, it can achieve both graph structure identification and parameters estimation simultaneously in high dimensional settings. The efficiency and accuracy are the main advantages of this novel method.
We provide finite sample theory and asymptotic theory for the extreme graphical lasso. In particular, we show a consistent identification of the graph and accurate estimation of the non-sparse parameters in $\Theta$. We provide practical guidance in choosing the tuning parameters for applying the extreme graphical lasso method, based on theoretical motivations. The performance of the proposed method is shown by simulations. Finally, we apply the extreme graphical lasso to two real data examples to illustrate its usefulness in uncovering underlying dependence structure of extreme events.
@@ 1.1 Notation
  • For matrix $A$ and vector $v$, $\|Av\|_{\infty}\leq|||A|||_{\infty}\|v\|_{\infty}$.
  • For matrices $A$ and $B$ with compatible dimensions, $|||AB|||_{\infty}\leq|||A|||_{\infty}|||B|||_{\infty}$.
@@ 2.1 Graphical models for extremes
where $\Lambda$ is a Radon measure on the cone $\mathcal{E}=[0,\infty)^{d}\backslash\{\mathbf{0}\}$ and $\Lambda(\mathbf{z})$ is shorthand for $\Lambda([0,\infty)^{d}\backslash[\mathbf{0},\mathbf{z}])$, the measure $\Lambda$ is known as the exponent measure and characterizes the dependence structure of $\mathbf{X}$ in the tail region, see, e.g. de Haan and Ferreira (2006, Chapter 6.1).
@@
Here the random vector $\mathbf{Y}$ is defined with support on the $L$-shaped set =$\mathcal{L}=\{\mathbf{x}\in\mathcal{E}:\|\mathbf{x}\|_{\infty}>1\}$. Its distribution is known as a multivariate Pareto distribution; see Rootzén and Tajvidi (2006).
The framework of graphical models for extremes (Engelke and Hitz, 2020) considers the conditional independence of the threshold exceedance limit $\mathbf{Y}$ in (2.2). Since $\mathbf{Y}$ is defined on the $L$-shaped set =$\mathcal{L}=\{\mathbf{x}\in\mathcal{E}:\|\mathbf{x}\|_{\infty}>1\}$, which is not a product space, the notion of conditional independence is instead defined on the subspace $\mathcal{L}^{k}=\{\mathbf{x}\in\mathcal{L}:x_{k}>1\}$ for each $k$. Define the random vector $\mathbf{Y}^{k}\stackrel{{\scriptstyle d}}{{=}}\mathbf{Y}|Y_{k}>1$. Then $\mathbf{Y}$ is said to exhibit conditional independence in extremes between component $i$ and $j$ if and only if $Y_{i}^{k}$ and $Y_{j}^{k}$ are conditionally independent for all $k$,
$Y_{i}\perp\!\!\!\perp_{e}Y_{j}|\mathbf{Y}_{\backslash\{i,j\}}\quad\Leftrightarrow\quad\forall k\in\{1,\ldots,d\}:Y_{i}^{k}\perp\!\!\!\perp Y_{j}^{k}|\mathbf{Y}^{k}_{\backslash\{i,j\}},$ (2.3)
where $\mathbf{Y}_{\backslash\{i,j\}}$ ($\mathbf{Y}^{k}_{\backslash\{i,j\}}$) indicates all other dimensions in $\mathbf{Y}$ ($\mathbf{Y}^{k}$) excluding $\{i,j\}$.
@@
$\{i,j\}\notin E\quad\Leftrightarrow\quad Y_{i}\perp\!\!\!\perp_{e}Y_{j}|\mathbf{Y}_{\backslash\{i,j\}}.$
@@
$Y_{i}\perp\!\!\!\perp_{e}Y_{j}|\mathbf{Y}_{\backslash\{i,j\}}\quad\Leftrightarrow\quad\lambda(\mathbf{y})=\frac{\lambda_{\backslash\{i\}}(\mathbf{y}_{\backslash\{i\}})\lambda_{\backslash\{j\}}(\mathbf{y}_{\backslash\{,j\}})}{\lambda_{\backslash\{i,j\}}(\mathbf{y}_{\backslash\{i,j\}})},$ (2.4)
where $\mathbf{y}_{\backslash{A}}$ denotes the entries of $\mathbf{y}$ outside of the index set $A$.
@@ 2.2.1 Charaterization using variagoram matrix $\Gamma$ and sub-covariance matrices $\Sigma^{(k)}$’s
Combining (2.5) and (2.6), there exists a one-to-one transformation between $\Gamma$ and $\tilde{\Sigma}^{(k)}$. Consequently, we can use $\tilde{\Sigma}^{(k)}$, for any fixed $k$, to parametrize an HR distribution. Note that $\tilde{\Sigma}^{(k)}$ is degenerate on the $k$-th row and the $k$-th column. Let $\Sigma^{(k)}$ be the $(d-1)\times(d-1)$ matrix constructed by removing the $k$-th row and the $k$-th column from $\tilde{\Sigma}^{(k)}$. For convenience, we index the rows and columns of $\Sigma^{(k)}$ using the original row and column numbers from $\tilde{\Sigma}^{(k)}$. By definition, $\tilde{\Sigma}^{(k)}$ is of rank $d-1$, therefore $\Sigma^{(k)}$ is of full rank.
@@
where $\phi_{d}(\cdot;\Sigma^{(k)})$ is the density of a centered $d$-dimensional Gaussian distribution with covariance matrix $\Sigma^{(k)}$ and $\mathbf{y}^{\prime}_{k}=\{\log(y_{i}/y_{k})+\Gamma_{ik}/2\}_{i=\{1,\ldots,d\}\backslash k}$. Note again that this formula holds for any $k=1,\ldots,d$.
@@ 2.2.2 Characterization using precision matrix $\Theta$
$Y_{i}\perp\!\!\!\perp_{e}Y_{j}|\mathbf{Y}_{\backslash\{i,j\}}\quad\Leftrightarrow\quad\Theta^{(k)}_{ij}=0,\quad\forall k\neq i,j.$
@@
$\Theta^{(k^{\prime})}_{ij}=\Theta^{(k)}_{ij}$ , if $\text{if }i,j\neq k;$
$\Theta^{(k^{\prime})}_{ik}=-\sum_{l\neq k}\Theta^{(k)}_{il}$ , if $\text{if }i\neq k\text{ and }j=k;$
$\Theta^{(k^{\prime})}_{kk}=\sum_{m,l\neq k}\Theta^{(k)}_{ml}$ , if $\text{if }i=k\text{ and }j=k.$
@@
$\Theta_{ij}=\Theta^{(k)}_{ij}$ , if $\text{if }i,j\neq k;$
$\Theta_{ik}=-\sum_{l\neq k}\Theta^{(k)}_{il}$ , if $\text{if }i\neq k\text{ and }j=k;$
$\Theta_{kk}=\sum_{m,l\neq k}\Theta^{(k)}_{ml}$ , if $\text{if }i=k\text{ and }j=k.$
@@
$Y_{i}\perp\!\!\!\perp_{e}Y_{j}|\mathbf{Y}_{\backslash\{i,j\}}\quad\Leftrightarrow\quad\Theta^{(k)}_{ij}=0,\,i,j\neq k\quad\Leftrightarrow\quad\Theta_{ij}=0.$
@@ 2.2.3 Characterization using covariance matrix $\Sigma$
In the following, we define a covariance matrix which serves as the ‘inverse’ fo $\Theta$. Given a variogram $\Gamma$, we consider the corresponding random vector $\mathbf{W}$ with the restriction $\sum_{k=1}^{d}W_{k}\equiv 0$.
@@
Let $\mathbf{W}$ be a random vector with variogram $\Gamma\in\mathcal{D}_{0}$. Then the random vector $\mathbf{W}^{\prime}:=\mathbf{W}-\bar{W}\cdot\mathbf{1}$ has variogram $\Gamma$ and covariance matrix
@@
=$\mathcal{L}=\{A\in\mathbb{R}^{d\times d}:A\succeq 0,A\mathbf{1}=\mathbf{0},\text{rank}(A)=d-1\}.$
@@ 3.1 Graphical lasso estimator of $\Theta$
Unlike from the classical graphical lasso, as $\gamma$ increases, this will not result in a sparse graph. Intuitively, due to the constraint that $\sum_{i=1}^{d}\Theta_{ij}=0$ for any $j$, $\Theta_{ij}$ will approach $-\Theta_{jj}/(d-1)$ as $\gamma$ increases, not 00. See, for example, Theorem 3.1 of Ying et al. (2020).
@@ 3.2 Solving the extreme graphical lasso
Denote $c:=\frac{1}{dM}$, $\Theta^{*}:=\Theta+\frac{1}{dM}\mathbf{1}\mathbf{1}$ and $S^{*}:=S+\frac{M}{d}\mathbf{1}\mathbf{1}^{T}$. Then the problem in (3.6) can be reformulated as
@@
This is similar to the Gaussian graphical lasso problem in (1.1) except that we penalize the off-diagonal entries of $\Theta^{*}$ to a positive constant $c$ instead of 00. Despite the similarity, this problem cannot be solved directly by the state-of-the-art GLASSO algorithm proposed in Friedman et al. (2008). Instead, we modify another block coordinate descent algorithm, the P-GLASSO algorithm, from Mazumder and Hastie (2012). Similar to GLASSO, P-GLASSO is computationally efficient and is guaranteed to converge on a trajectory of positive definite matrices. In Appendix E, we provide the algorithm for the extreme graphical lasso and its convergence properties.
@@ 4 Theoretical results
Given $M>0$, define $\Omega=\Sigma^{*}\otimes\Sigma^{*}$ where $\Sigma^{*}=\Sigma+\frac{M}{d}\mathbf{1}\mathbf{1}^{T}$ and $\otimes$ is the Kroneker product. We assume that there exists $10<\alpha<$ 1 such that
$|||\Omega_{E^{c}E}(\Omega_{EE})^{-1}|||_{\infty}<1-\alpha,$
@@
Assume that there exists $10<\epsilon<$ 1, such that
$\|S-\Sigma\|_{\infty}\leq\delta_{n}\leq\frac{(1-\epsilon)\epsilon\alpha^{2}}{D|||\Sigma^{*}|||_{\infty}|||\Omega_{EE}^{-1}|||_{\infty}\left[(1-\epsilon)\alpha+|||\Sigma^{*}|||_{\infty}^{2}|||\Omega_{EE}^{-1}|||_{\infty}\right]}.$
@@
$\overline{C}_{\gamma}$ ::=:= $\frac{(1-\epsilon)\alpha(1-\alpha)}{D|||\Sigma^{*}|||_{\infty}|||(\Omega_{EE})^{-1}|||_{\infty}\left[(1-\epsilon)\alpha+|||\Sigma^{*}|||_{\infty}^{2}|||(\Omega_{EE})^{-1}|||_{\infty}\right]},$ (4.1)
$\underline{C}_{\gamma}(\delta_{n})$ ::=:= $\frac{1-\alpha}{\epsilon\alpha}\cdot\delta_{n}.$ (4.2)
@@
$r_{n}:=\frac{|||\left(\Omega_{EE}\right)^{-1}|||_{\infty}}{1-\alpha}\cdot\gamma_{n}.$ (4.3)
@@
$P(\hat{E}=E)\to 1\text{\ \ and\ \ }\|\hat{\Theta}_{lasso}-\Theta\|_{\infty}=O_{P}\left(\delta_{n}\right).$
@@ 5.3 Choice of $M$: a theoretical guidance
The quantity $|||\Omega_{E^{c}E}(\Omega_{EE})^{-1}|||_{\infty}$ in Condition 4.1 can be bounded above by
$|||\Omega_{E^{c}E}(\Omega_{EE})^{-1}|||_{\infty}\leq\left(|E|\cdot\frac{\lambda_{\max}^{2}(\Omega)-\lambda_{\min}^{2}(\Omega)}{\lambda_{\min}^{2}(\Omega)}\right)^{1/2},$
@@
From Proposition 5.2, choosing $M$ smaller than $\lambda_{2}$ or larger than $\lambda_{d}$ will increase the upper bound for $|||\Omega_{E^{c}E}(\Omega_{EE})^{-1}|||_{\infty}$. Therefore we recommend the choice of $M\in[\lambda_{2},\lambda_{d}]$ to minimize this upper bound.
@@ 5.4 Choosing $M$: numerical examples
We now investigate two examples of $d=4$ and show how the choice of $M$ affects the mutual incoherence condition $|||\Omega_{E^{c}E}(\Omega_{EE})^{-1}|||_{\infty}$. Figure 1 shows the graph structures for the two examples, the star graph and the diamond graph. The Mutual Incoherence conditions for these two graphs in the classical graphical lasso setting are studied in Ravikumar et al. (2011).
[…]
@@ 5.4.1 Star graph
which reflects the star graph in Figure 1(a). The top two panels of Figure 2 show the values of $|||\Omega_{E^{c}E}(\Omega_{EE})^{-1}|||_{\infty}$ plotted against the values of $M$ for star graphs with parameters $x=0.8$ and $x=1.2$. As $[\lambda_{2},\lambda_{d}]$ is suggested as the practical range for $M$ from Section 5.3, the values of $\lambda_{2}$ and $\lambda_{d}$ are indicated as dash vertical lines in each graph. We observe that in both examples, $\lambda_{2}$ is a good indication of the lowest possible value of $|||\Omega_{E^{c}E}(\Omega_{EE})^{-1}|||_{\infty}$ while the majority of the interval $M\in[\lambda_{2},\lambda_{d}]$ satisfies the Mutual Incoherence condition.

### v3

@@ 1 Introduction
In this paper, we propose a direct estimate of $\Theta$ with a built-in option for sparse estimation via $L_{1}$-regularization. We term it the extreme graphical lasso (EGLasso). The core idea is as follows. We show that by adding a positive constant $c$ to each entry of $\Theta$, the matrix
@@
where $\gamma>0$ is the regularization parameter. The EGLasso thus achieves both graph structure identification and parameter estimation of $\Theta$ by solving a single penalized likelihood problem. Compared to the EGLearn procedure of Engelke et al. (2022), which requires solving $d$ separate optimization problems, EGLasso is computationally more efficient, particularly for large $d$. On the other hand, EGLearn with neighborhood selection can achieve better graph recovery for very sparse graphs, as it exploits the conditional independence structure node-by-node. The two approaches offer complementary strengths.
We provide finite sample theory and asymptotic theory for the extreme graphical lasso. In particular, we show a consistent identification of the graph and accurate estimation of the non-sparse parameters in $\Theta$. We provide practical guidance in choosing the tuning parameters for applying the extreme graphical lasso method, based on theoretical motivations. We compare the performance of EGLasso with EGLearn through an extensive simulation study using data simulated from the max-stable Hüsler-Reiss distribution. The simulation results show that EGLasso achieves comparable graph recovery for moderately dense graphs, while being 3355 times faster at dimension $d=100$. Finally, we apply EGLasso to two real data examples and compare the results with EGLearn.
@@ 1.1 Notation
[…]
@@ 2.1 Graphical models for extremes
[…]
@@
Here the random vector $\mathbf{Y}$ is defined with support on the $L$-shaped set $\mathcal{L}=\{\mathbf{x}\in\mathcal{E}:\|\mathbf{x}\|_{\infty}>1\}$. Its distribution is known as a multivariate Pareto distribution; see Rootzén and Tajvidi (2006).
The framework of graphical models for extremes (Engelke and Hitz, 2020) considers the conditional independence of the threshold exceedance limit $\mathbf{Y}$ in (2.2). Since $\mathbf{Y}$ is defined on the $L$-shaped set $\mathcal{L}=\{\mathbf{x}\in\mathcal{E}:\|\mathbf{x}\|_{\infty}>1\}$, which is not a product space, the notion of conditional independence is instead defined on the subspace $\mathcal{L}^{k}=\{\mathbf{x}\in\mathcal{L}:x_{k}>1\}$ for each $k$. Define the random vector $\mathbf{Y}^{k}\stackrel{{\scriptstyle d}}{{=}}\mathbf{Y}|Y_{k}>1$. Then $\mathbf{Y}$ is said to exhibit conditional independence in extremes between component $i$ and $j$ if and only if $Y_{i}^{k}$ and $Y_{j}^{k}$ are conditionally independent for all $k$,
[…]
@@
$i,j\}\notin E\quad\Leftrightarrow\quad Y_{i}\perp\!\!\!\perp_{e}Y_{j}|\mathbf{Y}_{\backslash\{i,j\}}.$
@@
[…]
@@ 2.2.1 Charaterization using variagoram matrix $\Gamma$ and sub-covariance matrices $\Sigma^{(k)}$’s
Combining (2.5) and (2.6), there exists a one-to-one transformation between $\Gamma$ and $\tilde{\Sigma}^{(k)}$. Consequently, we can use $\tilde{\Sigma}^{(k)}$, for any fixed $k$, to parametrize an HR distribution. Note that $\tilde{\Sigma}^{(k)}$ is degenerate on the $k$-th row and the $k$-th column. Let $\Sigma^{(k)}$ be the $(d-1)\times(d-1)$ matrix constructed by removing the $k$-th row and the $k$-th column from $\tilde{\Sigma}^{(k)}$. For convenience, we index the rows and columns of $\Sigma^{(k)}$ using the original row and column numbers from $\tilde{\Sigma}^{(k)}$. By definition, $\tilde{\Sigma}^{(k)}$ is of rank d1d-1, therefore $\Sigma^{(k)}$ is of full rank.
@@
[…]
@@ 2.2.2 Characterization using precision matrix $\Theta$
[…]
@@
$\Theta^{(k^{\prime})}_{ij}=\Theta^{(k)}_{ij}$ , $\text{if }i,j\neq k;$
$\Theta^{(k^{\prime})}_{ik}=-\sum_{l\neq k}\Theta^{(k)}_{il}$ , $\text{if }i\neq k\text{ and }j=k;$
$\Theta^{(k^{\prime})}_{kk}=\sum_{m,l\neq k}\Theta^{(k)}_{ml}$ , $\text{if }i=k\text{ and }j=k.$
@@
$\Theta_{ij}=\Theta^{(k)}_{ij}$ , $\text{if }i,j\neq k;$
$\Theta_{ik}=-\sum_{l\neq k}\Theta^{(k)}_{il}$ , $\text{if }i\neq k\text{ and }j=k;$
$\Theta_{kk}=\sum_{m,l\neq k}\Theta^{(k)}_{ml}$ , $\text{if }i=k\text{ and }j=k.$
@@
[…]
@@ 2.2.3 Characterization using covariance matrix $\Sigma$
In the following, we define a covariance matrix which serves as the ‘inverse’ fo $\Theta$. Given a variogram $\Gamma$, we consider the corresponding random vector $\mathbf{W}$ with the restriction $\sum_{k=1}^{d}W_{k}\equiv 0$. The following proposition shows that, given any variogram $\Gamma\in\mathcal{D}_{0}$, one can always construct a random vector whose components sum to zero and whose variogram equals $\Gamma$, and that such a vector admits a well-defined covariance matrix.
@@
Let $\mathbf{W}$ be a random vector with variogram $\Gamma\in\mathcal{D}_{0}$. Then the random vector $\mathbf{W}^{\prime}:=\mathbf{W}-\bar{W}\cdot\mathbf{1}$, which satisfies $\sum_{k=1}^{d}W^{\prime}_{k}=0$, has variogram $\Gamma$ and covariance matrix
@@
$\mathcal{L}=\{A\in\mathbb{R}^{d\times d}:A\succeq 0,A\mathbf{1}=\mathbf{0},\text{rank}(A)=d-1\}.$
@@ 3.1 Graphical lasso estimator of $\Theta$
Unlike from the classical graphical lasso, as $\gamma$ increases, this will not result in a sparse graph. Intuitively, due to the constraint that $\sum_{i=1}^{d}\Theta_{ij}=0$ for any $j$, $\Theta_{ij}$ will approach $\Theta_{jj}/(d-1)$ as $\gamma$ increases, not 0. See, for example, Theorem 3.1 of Ying et al. (2020).
@@ 3.2 Solving the extreme graphical lasso
Denote $c:=\frac{1}{dM}$, $\Theta^{*}:=\Theta+\frac{1}{dM}\mathbf{1}\mathbf{1}^{T}$ and $S^{*}:=S+\frac{M}{d}\mathbf{1}\mathbf{1}^{T}$. Then the problem in (3.6) can be reformulated as
@@
This is similar to the Gaussian graphical lasso problem in (1.1) except that we penalize the off-diagonal entries of $\Theta^{*}$ to a positive constant $c$ instead of 0. Despite the similarity, this problem cannot be solved directly by the state-of-the-art GLASSO algorithm proposed in Friedman et al. (2008). Instead, we modify another block coordinate descent algorithm, the P-GLASSO algorithm, from Mazumder and Hastie (2012). Similar to GLASSO, P-GLASSO is computationally efficient and is guaranteed to converge on a trajectory of positive definite matrices. In Appendix E, we provide the algorithm for the extreme graphical lasso and its convergence properties.
@@ 4 Theoretical results
Given $M>0$, define $\Omega=\Sigma^{*}\otimes\Sigma^{*}$ where $\Sigma^{*}=\Sigma+\frac{M}{d}\mathbf{1}\mathbf{1}^{T}$ and $\otimes$ is the Kroneker product. We assume that there exists $0<\alpha<1$ such that
$|\Omega_{E^{c}E}(\Omega_{EE})^{-1}|||_{\infty}<1-\alpha,$
@@
Assume that there exists $0<\epsilon<1$, such that
[…]
@@
[…]
$\underline{C}_{\gamma}(\delta_{n})$ $:=$ $\frac{1-\alpha}{\epsilon\alpha}\cdot\delta_{n}.$ (4.2)
@@
[…]
@@
In particular, if $\min\{|\Theta_{ij}|;(i,j)\in E,i\neq j\}>r_{n}$, then $\hat{E}=E$.
Remark 4.2.
The condition $\min\{|\Theta_{ij}|;(i,j)\in E,i\neq j\}>r_{n}$ is a minimal signal strength requirement: true edges can only be recovered if the corresponding entries in $\Theta$ are sufficiently large relative to the estimation error $r_{n}$. This condition is standard in the graphical lasso literature; see, e.g., Theorem 1 in Ravikumar et al. (2011). In Theorem 4.3 below, this condition is absorbed into the requirement that $\delta_{n}^{-1}\min\{|\Theta_{ij}|;(i,j)\in E\}\to\infty$.
@@
[…]
@@ 5.3 Choice of $M$: a theoretical guidance
The quantity $|\Omega_{E^{c}E}(\Omega_{EE})^{-1}|||_{\infty}$ in Condition 4.1 can be bounded above by
$|\Omega_{E^{c}E}(\Omega_{EE})^{-1}|||_{\infty}\leq\left(|E|\cdot\frac{\lambda_{\max}^{2}(\Omega)-\lambda_{\min}^{2}(\Omega)}{\lambda_{\min}^{2}(\Omega)}\right)^{1/2},$
@@
From Proposition 5.2, choosing $M$ smaller than $\lambda_{2}$ or larger than $\lambda_{d}$ will increase the upper bound for $|\Omega_{E^{c}E}(\Omega_{EE})^{-1}|||_{\infty}$. Therefore we recommend the choice of $M\in[\lambda_{2},\lambda_{d}]$ to minimize this upper bound.
@@ 5.4 Choosing $M$: numerical examples
We now investigate two examples of $d=4$ and show how the choice of $M$ affects the mutual incoherence condition $|\Omega_{E^{c}E}(\Omega_{EE})^{-1}|||_{\infty}$. Figure 1 shows the graph structures for the two examples, the star graph and the diamond graph. The Mutual Incoherence conditions for these two graphs in the classical graphical lasso setting are studied in Ravikumar et al. (2011).
$X_{1}$$X_{2}$$X_{3}$$X_{4}$$X_{2}$$X_{1}$$X_{4}$$X_{3}$(a)(b)
@@ 5.4.1 Star graph
which reflects the star graph in Figure 1(a). The top two panels of Figure 2 show the values of $|\Omega_{E^{c}E}(\Omega_{EE})^{-1}|||_{\infty}$ plotted against the values of $M$ for star graphs with parameters $x=0.8$ and $x=1.2$. As $\lambda_{2},\lambda_{d}]$ is suggested as the practical range for $M$ from Section 5.3, the values of $\lambda_{2}$ and $\lambda_{d}$ are indicated as dash vertical lines in each graph. We observe that in both examples, $\lambda_{2}$ is a good indication of the lowest possible value of $|\Omega_{E^{c}E}(\Omega_{EE})^{-1}|||_{\infty}$ while the majority of the interval $M\in[\lambda_{2},\lambda_{d}]$ satisfies the Mutual Incoherence condition.